# Picasso: Memory-Efficient Graph Coloring Using Palettes With Applications in Quantum Computing

### 12 Jan 2024 v1

@@ I Introduction
Given a graph $G(V,E)$, the problem of graph coloring is to assign a color to each vertex such that no two adjacent vertices are assigned the same color, while minimizing the number of colors used. Graph coloring is one of the central problems in combinatorial optimization and it has applications in various scientific domains [1]. Many algorithmic techniques have been developed to solve coloring in sequential, parallel, and distributed settings for many variants of the coloring problem [2]. Most of them require loading the entire graph and several auxiliary data structures into memory and, therefore, they are avid of memory. For massive graphs, especially on limited-memory accelerators (GPUs), these algorithms easily exhaust the available memory, and the problem is exacerbated for dense graphs due to quadratic scaling of edges. Our work considers graphs that are $\approx 50\%$ dense ($|E|\approx|V|^{2}/2$) for which the current graph coloring approaches [3, 4, 5, 6] quickly run out of memory when processing large instances (§ VII).
@@
One near-term solution to improve the scaling of quantum algorithms is the unitary partitioning and its variants, which aim to find compact representations of a linear combination of Pauli strings (§ II). These strategies, predominantly based on the graph analysis of Pauli strings, typically yield a reduction ranging from 1/10 to 1/6 in the number of alternative unitaries for small test cases. The unitary partitioning problem boils down to a graph coloring problem, where the vertices of the graph represent the Pauli strings, and the edges represent whether these Pauli strings obey a mathematical relation (§ II). These graphs are large and dense. Therefore, the state-of-the-art approaches for the unitary partitioning problem can only solve small molecules with a few thousand Pauli strings, whereas the desired scale is on the order of $\mathcal{O}(10^{6}\sim 10^{12})$ of Pauli strings.
Building on recent developments in sublinear algorithms for graph coloring [10], we introduce a novel parameterized coloring algorithm, Picasso, where the parameters of the algorithm provide a trade-off between the quality of the solution (i.e., the number of colors used) and the resource consumption of the algorithm. We theoretically prove that Picasso has a sublinear space requirement (§ IV), which also makes it attractive to implement on GPUs. We empirically show that Picasso can solve trillion edge graph problems arising from molecules with more than 2 million Pauli strings in under fifteen minutes. Many of the results we report are the first-of-its-kind for the corresponding molecule and basis set. Since our parametric algorithm considers the target number of colors, we show that high-quality coloring can be achieved by an aggressive choice of parameters. We design a machine learning prediction model to determine the parameters configuration to be used to achieve a given trade-off between coloring quality and resource requirements (§VI).
@@ III Related Work
The recent seminal work of Assadi, Chen and Khanna [10] (ACK, henceforth), stduied the $(\Delta+1)$-coloring problem with sublinear space and time constraints. They developed the Pallete Sparsification Theorem, which reduces the $\Delta+1$-coloring of a graph $G(V,E)$ to a list-coloring problem in a subgraph of $G$ with only $\mathcal{O}(|V|\log^{2}{|V|})$ edges. Applications of this theorem is shown by designing algorithms in dynamic semi-streaming in a single pass, sublinear query and MPC models. In terms of semi-streaming algoirthm, the only known $(\Delta+1)$-coloring before ACK’s algorithm [10] was the $\mathcal{O}(\log{|V|})$ pass distributed algorithm of Luby [21, 28], simulated in streaming setup. In this paper, we put ACK’s algorithm in practice by a few modification which is discussed in the following section.
IV Our Algorithm
Picasso attempts to color the graph by initially assigning to each vertex a list of candidate colors chosen uniformly at random from a palette of $\mathcal{P}$ colors $\{0,1,\ldots,\mathcal{P}-1\}$, and progressively assigning to each vertex a color from its color list without violating the graph coloring constraint. Picasso is inspired by the seminal ACK’s streaming algorithm [10]. By choosing the size of the palette as $\Delta+1$, the size of the color list as $\mathcal{O}(\log|V|)$, and adopting a special post-processing step, ACK’s algorithm can color a graph with maximum degree $\Delta$ with $\Delta+1$ colors in a single pass in sublinear space and time. We extend and modify this algorithm in several ways:
@@ IV Our Algorithm
Picasso, listed in Algorithm 1, takes a graph $G$ as input and computes a valid coloring of $G$ in the $color$ array. The algorithm starts with the original graph $G$, and iteratively computes coloring for a subgraph. In each iteration ($\ell$), it estimates the palette size ($\mathcal{P}_{\ell}$) and list size ($\mathcal{L}_{\ell})$ for the current subgraph $G_{\ell}$ (Line 6). Next, it assigns to each vertex a list of candidate colors chosen uniformly at random from the palette (Line 7). For each vertex, we record the list of colors in the $colList$ data structure, where $colList(u)$ refers to the list of colors assigned to vertex $u$. We build the conflict subgraph, which consists of the edges that share a common color in their corresponding lists (Line 8) using the $colList$ array. We then color the unconflicted vertices using an arbitrary color from their color list (Line 9), and attempt to color the conflict graph (Line 10). The vertices uncolored in the current iteration are denoted as $V_{u}$. We then compute the subgraph induced by $V_{u}$ and continue if $V_{u}$ is non-empty. We note that the algorithm in every iteration starts with a new palette of colors and attempts to color the subgraph with these new colors. The colors of an iteration are not reused in the subsequent iterations. We can ensure that by defining the palette set as $\{(\ell-1)\mathcal{P},\ldots\ell\mathcal{P}\}$ at iteration $\ell$.
@@ IV-B Coloring the Conflict Graph
Static order schemes: Given an ordered set of vertices, we iterate through the set and attempt to color each vertex with the first available color in its list that does not conflict with the colors already assigned to its neighbors. For ordering, we can use Natural, Largest Degree First, Smallest Degree Last, or Random ordering, which have been explored for graph coloring [2].
Dynamic vertex order scheme: A second approach is to color the conflict graph using the lists in a dynamic order [29]. Here, we describe a dynamic greedy algorithm from [29] for the list coloring and discuss an efficient implementation of it. The algorithm attempts to color the vertices that are most constrained, i.e., it colors a vertex according to the size of the list. The vertex with the list with the smallest size is colored first; removing this color from the neighboring vertices changes the ordering dynamically. A naïve implementation of the dynamic list coloring algorithm would require $\mathcal{O}(|V_{c}|^{2}+|E_{c}|\mathcal{L})$ time. We have at most $|V_{c}|$ iterations, and in each iteration, we need to find the vertex with the smallest current list size in $\mathcal{O}(|V_{c}|)$ time, then color the vertex and mark this color removed from all its neighbors in $\mathcal{O}(\delta(v)\mathcal{L})$ time, where $\delta(v)$ is the degree of vertex $v$. Summing over $|V|$ gives us the total time. We can reduce it to $\mathcal{O}((|V_{c}|+|E_{c}|\mathcal{L})\log{|V_{c}|})$ using a minimum heap data structure with logarithmic update_key operation. Next, we present an efficient implementation in Algorithm 2 that eliminates the $\log{|V_{c}|}$ factor by using the bucketing technique.
@@
1:Input: Conflict graph $G_{c}$, list of colors $colList(v)$, $\forall v\in V_{c}$.
@@
Algorithm 2 stores the vertices of $G_{c}$ in an array of buckets $B$, according to the size of their color lists. The bucket at $B[i]$ holds the vertices $v\in V$, whose size of the $colList(v)$ is $i$. We define the lowest bucket from the array as the non-empty bucket with the smallest index. The algorithm marks all the vertices as unprocessed and continue until all the vertices are processed. In each iteration, the algorithm finds the lowest bucket and chooses a vertex uniformly at random from it. This vertex is colored with an arbitrary color from its list and marked as processed. We say the color is $c$. Then, the algorithm scans all neighbors of this vertex in $G_{c}$ and removes $c$ if it exists in their color list. If any neighbor’s list becomes empty, we mark this neighboring vertex as processed. The runtime of Algorithm 2 is $\mathcal{O}((|V_{c}|+|E_{c}|)\mathcal{L})$, since there might be at most $|V_{c}|$ iterations and in each iteration we require $\mathcal{O}(\mathcal{L})$ time to find the lowest bucket. The removals of a vertex from a bucket (Lines 9 and 14) can be implemented in constant time by an auxiliary array that stores the location of the vertex in the bucket, while the removal of a color from $colList$ (Line 10) takes $\mathcal{O}(\mathcal{L})$ Thus, processing a chosen vertex requires $\mathcal{O}(\delta(v)\mathcal{L})$ time, and summation over all vertices gives the total time.
@@ V Parallel GPU implementation
1:Input: Set of Pauli strings, $V$ and their list of colors $colList$
@@
The conflict subgraph construction on Line 8 of Algorithm 1 significantly dominates the execution time for our application. Because of this, we implemented a parallel GPU version to reduce the bottleneck. We note that, despite the original graph being dense, in almost all the cases the conflict graph is expected to be significantly sparse (details in §VII-A1). Although the pairwise comparisons of vertices are independent of each other, due to the unknown number of conflicting edges at runtime, we designed an implementation capable of generating the conflict graph in Compressed Sparse Row (CSR) format that is not only efficient in memory usage during construction but also enables contiguous access of memory for processing the conflict graph. We present this implementation in Algorithm 3. We first copy the input data on the GPU with a size of $N\mathcal{L}|V|/10$ 4-byte values for the encoded Pauli strings and assigned colors, and assign $2|V|$ 4-byte or 8-byte edge offset counters. If $|V|^{2}\geq 2^{32}$, then 8-byte offset counters will be used; otherwise, the offset counters are 4-bytes. All remaining available memory on the GPU, or the worst-case edgelist size of $2|V|(|V|-1)$, whichever is smaller, is then allocated to store the unordered edgelist, and the conflict graph generation kernel is launched on the GPU. For this kernel, each thread processes one of the $|V|(|V|-1)/2$ possible edges and, if it calculates that the edge is both a complement and conflicting edge, the edge is written in the output and the respective edge offsets are incremented. After the kernel execution, we are left with an unordered edge list of size $|E_{c}|$, assuming that each unordered edge is represented only once. Note that, since the edges of the complement graph are determined dynamically during conflict graph construction, the complement graph does not need to be stored on the GPU memory. If $|E_{c}|$ used less than half of the available GPU memory, then we generate the CSR output on the GPU. Otherwise, we read the unordered edge list and convert it to CSR on the host CPU instead. We attempted warp-level reduction on the offset values to condense the number of global atomic operations, but the number of conflicting edges was much smaller than the total number of possible edges ($\leq$5in most cases), so the overhead outweighed the benefits. In order to get a good estimation for $|E_{c}|$ for memory allocations, we developed a machine learning based predictor, which we define next (§VI).
@@ VII Experimental Evaluation
We evaluate the performance and quality of the solution obtained using Picasso to the current state-of-the-art approaches. For our evaluation, we used a machine equipped with a 64-core AMD EPYC 7742 CPU, 512GB host DDR memory, and an NVIDIA A100 GPU with 40GB of HBM memory. Table II lists the datasets we selected for our evaluation. These systems were chosen with a deliberate intent to encompass a broad spectrum of quantum scenarios. By picking $H_{n}$ molecular systems with varying values of $n=4,6,8,10$, we aimed to ensure diversity in system size, ranging from simple to complex structures, enabling us to gauge the algorithm’s scalability and performance across different magnitudes. The incorporation of three spatial configurations for each $H_{n}$ molecular system, namely 1D, 2D, and 3D, was intended to introduce dimensional variability, shedding light on the tool’s capability to manage problems with different geometric complexities and symmetries. As the size of the $H_{n}$ system increases, so does the intricacy of electron-electron interactions and correlations. By integrating systems of different sizes, the goal was to critically assess the tool’s proficiency in addressing varying degrees of electron correlation, a pivotal aspect in quantum calculations. Moreover, this diverse selection, spanning across multiple system sizes and dimensions, serves as an effective stress test for Picasso. It not only offers insights into its performance benchmarks but also aids in pinpointing potential areas of enhancement, ensuring a comprehensive evaluation of its reliability and robustness. We classify our datasets into three categories: i) Small ($\leq 10$ Billion edges); ii) Medium ($\leq 1$ Trillion edges); and iii) Large ($>1$ Trillion edges). We evaluate our implementation by considering:
@@ VII-C Performance Comparison with Kokkos-EB and ECL-GC-R
Figure 4: Relative performance comparison of $\mathcal{P}$ and Kokkos-EB on medium datasets size normalized to ECL-GC-R execution time. For Picasso runs, $\alpha=4.5$
@@ VII-D Parameter Sensitivity
The heatmap in Fig. 5 shows the impact of $\mathcal{P}$ and $\alpha$ on the final colors, conflicting edges, and execution time for a representative input of the general trend (H4_2D_6311g). The heatmaps show normalized quantities with respect to what is observed for the same input graph. Therefore, as a measure of quality we report the fraction of colors with respect to the number of vertices in the input (lower is better) while we report execution time and the normalized fraction of the total number of conflicting edges processed by the algorithm as measures of the work done by the algorithm (lower is better).

### 12 Feb 2024 v2

@@ I Introduction
Given a graph $G(V,E)$, the problem of graph coloring is to assign a color to each vertex such that no two adjacent vertices are assigned the same color, while minimizing the number of colors used. Graph coloring is one of the central problems in combinatorial optimization with applications in various scientific domains [1]. Many algorithmic techniques have been developed to solve coloring in sequential, parallel, and distributed settings for many variants of the coloring problem [2]. All of these algorithms are memory-demanding since they require loading the entire graph and several auxiliary data structrues into the memory. For massive graphs, especially on limited-memory accelerators (GPUs), these algorithms easily exhaust the available memory, and the problem is exacerbated for dense graphs due to quadratic scaling of edges. Our work considers graphs that are $\approx 50\%$ dense ($|E|\approx|V|^{2}/2$) for which the current graph coloring approaches [3, 4, 5, 6] quickly run out of memory when processing large instances (§ VII).
@@
One near-term solution to improve the scaling of quantum algorithms is the unitary partitioning and its variants, which aim to find compact representations of a linear combination of Pauli strings (§ II). These strategies, predominantly based on the graph analysis of Pauli strings, typically yield a reduction ranging from 1/10 to 1/6 in the number of alternative unitaries for small test cases. The unitary partitioning problem reduces to a graph coloring problem, where the vertices of the graph represent the Pauli strings, and the edges represent whether these strings obey an anticommute relation (§ II). These graphs are large and dense. Therefore, the state-of-the-art approaches for the unitary partitioning problem can only solve small molecules with a few thousand Pauli strings, whereas the desired scale is on the order of $\mathcal{O}(10^{6}\sim 10^{12})$ of Pauli strings.
Building on recent developments in sublinear algorithms for graph coloring [10], we introduce a novel parameterized coloring algorithm, Picasso, where the parameters of the algorithm provide a trade-off between the quality of the solution (i.e., the number of colors used) and the resource consumption of the algorithm. We theoretically prove that Picasso has a sublinear space requirement (§ IV), making it attractive to implement on GPUs. We empirically show that Picasso can solve trillion edge graph problems arising from molecules with more than 2 million Pauli strings in under fifteen minutes. Many of the results we report are the first-of-its-kind for the corresponding molecule and basis set. Since our parametric algorithm considers the target number of colors, we show that high-quality coloring can be achieved by an aggressive choice of parameters (§ VII-A1). We design a machine learning prediction model to determine the parameters configuration to be used to achieve a given trade-off between coloring quality and resource requirements (§VI).
@@ III Related Work
Graph Coloring in sublinear space: The recent seminal work of Assadi, Chen and Khanna [10] (ACK, henceforth), stduied the $(\Delta+1)$-coloring problem with sublinear space and time constraints. They developed the Pallete Sparsification Theorem, which reduces the $\Delta+1$-coloring of a graph $G(V,E)$ to a list-coloring problem in a subgraph of $G$ with only $\mathcal{O}(|V|\log^{2}{|V|})$ edges. Applications of this theorem is shown by designing algorithms in dynamic semi-streaming in a single pass, sublinear query and MPC models. In terms of semi-streaming algoirthm, the only known $(\Delta+1)$-coloring before ACK’s algorithm [10] was the $\mathcal{O}(\log{|V|})$ pass distributed algorithm of Luby [21, 28], simulated in streaming setup. In semi-streaming model, ACK uses palette size as $\Delta+1$, list size of colors as $\mathcal{O}(\log|V|)$ for each vertex, and a special post-processing step. In this paper, we put ACK’s algorithm in practice by non-trivial modifications listed as follows.
@@ IV Our Algorithm
Picasso, listed in Algorithm 1, attempts to color the graph by initially assigning to each vertex a list of candidate colors chosen uniformly at random from a palette of $\mathcal{P}$ colors $\{0,1,\ldots,\mathcal{P}-1\}$, and progressively assigning to each vertex a color from its color list without violating the graph coloring constraint. It takes a graph $G$ as input and computes a valid coloring of $G$ in the $color$ array. The algorithm starts with the original graph $G$, and iteratively computes coloring for a subgraph. In each iteration ($\ell$), it estimates the palette size ($\mathcal{P}_{\ell}$) and list size ($\mathcal{L}_{\ell})$ for the current subgraph $G_{\ell}$ (Line 6). Next, it assigns to each vertex a list of candidate colors chosen uniformly at random from the palette (Line 7). For each vertex, we record the list of colors in the $colList$ data structure, where $colList(u)$ refers to the list of colors assigned to vertex $u$. We build the conflict subgraph, which consists of the edges that share a common color in their corresponding lists (Line 8) using the $colList$ array. We then color the unconflicted vertices using an arbitrary color from their color list (Line 9), and attempt to color the conflict graph (Line 10). The vertices uncolored in the current iteration are denoted as $V_{u}$. We then compute the subgraph induced by $V_{u}$ and continue if $V_{u}$ is non-empty. We note that the algorithm in every iteration starts with a new palette of colors and attempts to color the subgraph with these new colors. The colors of an iteration are not reused in the subsequent iterations. We can ensure that by defining the palette set as $\{(\ell-1)\mathcal{P},\ldots\ell\mathcal{P}\}$ at iteration $\ell$.
@@ IV-B Coloring the Conflict Graph
Static order schemes: Given an ordered set of vertices, we iterate through the set and attempt to color each vertex with the first available color in its list that does not conflict with the colors already assigned to its neighbors. We may use popular [2] vertex order such as the Natural, Largest Degree First, Smallest Degree Last, or Random ordering.
Dynamic vertex order scheme: A second approach is to color the conflict graph using the lists in a dynamic order [29]. Here, we describe a dynamic greedy algorithm from [29] for the list coloring and discuss an efficient implementation of it. The algorithm attempts to color the vertices that are most constrained, i.e., it colors a vertex according to the size of the list. When a vertex is assigned a color from its list, the assigned color is removed from the list of all the neighbor of the vertex, rendering a dynamic order on the vertices. A naïve implementation of the dynamic list coloring algorithm would require $\mathcal{O}(|V_{c}|^{2}+|E_{c}|\mathcal{L})$ time. We have at most $|V_{c}|$ iterations, and in each iteration, we need to find the vertex with the smallest current list size in $\mathcal{O}(|V_{c}|)$ time, then color the vertex and mark this color removed from all its neighbors in $\mathcal{O}(\delta(v)\mathcal{L})$ time, where $\delta(v)$ is the degree of vertex $v$. Summing over $|V|$ gives us the total time. We can reduce it to $\mathcal{O}((|V_{c}|+|E_{c}|\mathcal{L})\log{|V_{c}|})$ using a minimum heap data structure with logarithmic update_key operation. Next, we present an efficient implementation in Algorithm 2 that eliminates the $\log{|V_{c}|}$ factor by using the bucketing technique.
@@
1:Input: Conflict graph $G_{c}$, color list $colList(v)$, $\forall v\in V_{c}$.
@@
Algorithm 2 stores the vertices of $G_{c}$ in an array of buckets $B$, according to the size of their color lists. The bucket at $B[i]$ holds the vertices $v\in V$, whose size of the $colList(v)$ is $i$. We define the lowest bucket from the array as the non-empty bucket with the smallest index. The algorithm marks all the vertices as unprocessed and continue until all the vertices are processed. In each iteration, the algorithm finds the lowest bucket and chooses a vertex uniformly at random from it. This vertex is colored with an arbitrary color from its list and marked as processed. We say the color is $c$. Then, the algorithm scans all neighbors of this vertex in $G_{c}$ and removes $c$ if it exists in their color list. If any neighbor’s list becomes empty, we mark this neighboring vertex as processed. The runtime of Algorithm 2 is $\mathcal{O}((|V_{c}|+|E_{c}|)\mathcal{L})$, since there might be at most $|V_{c}|$ iterations and in each iteration we require $\mathcal{O}(\mathcal{L})$ time to find the lowest bucket. The removals of a vertex from a bucket (Lines 9 and 14) can be implemented in constant time by an auxiliary array that stores the location of the vertex in the bucket, while the removal of a color from $colList$ (Line 10) takes $\mathcal{O}(\mathcal{L})$. Thus, processing a chosen vertex requires $\mathcal{O}(\delta(v)\mathcal{L})$ time, and summation over all vertices gives the total time.
@@ V Parallel GPU implementation
1:Input: Pauli strings, $V$ and their list of colors $colList$
@@
The conflict subgraph construction on Line 8 of Algorithm 1 significantly dominates the execution time for our application. Because of this, we implemented a parallel GPU version to reduce the bottleneck. We note that, despite the original graph being dense, in almost all the cases the conflict graph is expected to be significantly sparse (details in §VII-A1). Although the pairwise comparisons of vertices are independent of each other, due to the unknown number of conflicting edges at runtime, we designed an implementation capable of generating the conflict graph in Compressed Sparse Row (CSR) format that is not only efficient in memory usage during construction but also enables contiguous access of memory for processing the conflict graph. We present this implementation in Algorithm 3. As a preprocessing (now shown in the Algorithm 3), we copy the input data on the GPU with a size of $N\mathcal{L}|V|/10$ 4-byte values for the encoded Pauli strings and their list of colors, and initialize $2|V|$ edge offset counters. We use 8 bytes for the counter if $|V|^{2}\geq 2^{32}$; otherwise, the offset counters are 4-bytes. All remaining available memory on the GPU, or the worst-case edgelist size of $2|V|(|V|-1)$, whichever is smaller, is then allocated to store the unordered edgelist, and the conflict graph generation kernel is launched on the GPU. For this kernel, each thread processes one of the $|V|(|V|-1)/2$ possible edges and, and inspect whether the edge is both a complement and conflicting. If so, the edge is written in the output and the respective edge offsets are incremented. After the kernel execution, we are left with an unordered edge list of size $|E_{c}|$.The edges of the complement graph are determined independently during conflict graph construction, and the complement graph does not need to be stored on the GPU memory. Since for CSR representation each edge is stored twice, if $|E_{c}|$ used less than half of the available GPU memory, then we generate the CSR output on the GPU. Otherwise, we read the unordered edge list and convert it to CSR on the host CPU instead. We attempted warp-level reduction on the offset values to condense the number of global atomic operations, but the number of conflicting edges was much smaller than the total number of possible edges ($\leq$5in most cases), so the overhead outweighed the benefits. To get an estimation for $|E_{c}|$ to preallocate memory on the GPU, we developed a machine learning based predictor, which we describe next (§VI).
@@ VII Experimental Evaluation
We evaluate the performance and quality of the solution obtained using Picasso to the current state-of-the-art approaches. For our evaluation, we used a machine equipped with a 64-core AMD EPYC 7742 CPU, 512GB host DDR memory, and an NVIDIA A100 GPU with 40GB of HBM memory. Table II lists the datasets we selected for our evaluation. These systems are chosen with a deliberate intent to encompass a broad spectrum of quantum scenarios. By picking $H_{n}$ molecular systems with varying values of $n=4,6,8,10$, we aim to ensure diversity in system size, ranging from simple to complex structures, enabling us to gauge the algorithm’s scalability and performance across different magnitudes. The incorporation of three spatial configurations for each $H_{n}$ molecular system, namely 1D, 2D, and 3D, introduces dimensional variability, shedding light on the tool’s capability to manage problems with different geometric complexities and symmetries. As the size of the $H_{n}$ system increases, so does the intricacy of electron-electron interactions and correlations. By integrating systems of different sizes, the goal is to critically assess the tool’s proficiency in addressing varying degrees of electron correlation, a pivotal aspect in quantum calculations. Moreover, this diverse selection, spanning across multiple system sizes and dimensions, serves as an effective stress test for Picasso. It not only offers insights into its performance benchmarks but also aids in pinpointing potential areas of enhancement, ensuring a comprehensive evaluation of its reliability and robustness. We classify our datasets into three categories: i) Small ($\leq 10$ Billion edges); ii) Medium ($\leq 1$ Trillion edges); and iii) Large ($>1$ Trillion edges). We evaluate our implementation by considering:
@@ VII-C Performance Comparison with Kokkos-EB and ECL-GC-R
Figure 4: Performance comparison of Picasso and Kokkos-EB on the small datasets relative to ECL-GC-R execution time. For Picasso runs, $\mathcal{P}$ is varied and $\alpha=4.5$
@@ VII-D Parameter Sensitivity
The heatmap in Fig. 5 shows the impact of $\mathcal{P}$ and $\alpha$ on the final colors, conflicting edges, and execution time for a representative input (H4_2D_6311g). The heatmaps show normalized quantities with respect to what is observed for the same input graph. Therefore, as a measure of quality we report the fraction of colors with respect to the number of vertices in the input (lower is better) while we report execution time and the normalized fraction of the total number of conflicting edges processed by the algorithm as measures of the work done by the algorithm (lower is better).